OpenStudy (anonymous):

What is S5 for 1000 + 500 + 250 + ... ?

OpenStudy (anonymous):

@stamp @PeterPan

OpenStudy (anonymous):

This series is not like the series we had earlier, which was an arithmetic series... This is a new and more cunning type of series o.O

OpenStudy (stamp):

It is a pattern. What is the pattern? Think different.

OpenStudy (anonymous):

I'm trying to figure that pattern out

OpenStudy (stamp):

@Kayy_Drizzyy Take your time.

OpenStudy (stamp):

@Kayy_Drizzyy Hint: It is geometric. If you do not know what geometric sequence is, please open your textbook to the index, find the "geometric sequences" page number, and read the section.

OpenStudy (anonymous):

Idk what the patterm is

OpenStudy (anonymous):

is it \(S_5\) or is it \[\sum_{n=0}^{\infty}1000\left(\frac{1}{2}\right)^n\]?

OpenStudy (anonymous):

So my pattern is (1/2) ? right because i did 1000(1/2) and got my next answer which was 500

OpenStudy (anonymous):

o.O @Kayy_Drizzyy Don't give up :) Look at the second term, divide it by the first. Look at the third term, divide it by the second. Notice a pattern? :)

OpenStudy (anonymous):

Oh you have it :)

OpenStudy (anonymous):

The problem seems to go down not you means 100(1/2)=500 500(1/2)=250 250(1/2)=125 125(1/2)=62.5

OpenStudy (anonymous):

My answer problem are large numbers . How do i get that answer

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

add

OpenStudy (anonymous):

add by (1/2) ?

OpenStudy (anonymous):

This is what is known as a Geometric Series.... Each term is just the previous term, multiplied to a constant known as... ... the COMMON RATIO

OpenStudy (anonymous):

\(S_5\) if i read it correctly means you have exactly 5 terms. Add up the first five numbers

OpenStudy (anonymous):

i know that but what am i adding by to get my S5 term ?

OpenStudy (anonymous):

Which as you pointed out, is 1/2 Luckily, a formula exists for getting the sum of the first n-terms of a geometric series :) \[\huge S_n = a\left( \frac{1-r^n}{1-r} \right)\] Where as is your first term and r is your... ... COMMON RATIO >:D

OpenStudy (anonymous):

I'm confused now

OpenStudy (anonymous):

Well, what's your common ratio r? It's (1/2) right?

OpenStudy (anonymous):

yes r is 1/2 but where does the A(1-1/2n) come from to get my answer ?

OpenStudy (anonymous):

a is the very first term, which happens to be 1000 :)

OpenStudy (anonymous):

and n is the number of terms you're adding, which happens to be 5.

OpenStudy (anonymous):

okay so break it down to me so i can solve i it and get my answer.

OpenStudy (anonymous):

\[\huge S_n = a\left( \frac{1-r^n}{1-r} \right)\] You're taking the sum of the first 5 terms, so n=5

OpenStudy (anonymous):

\[\huge S_5 = a\left( \frac{1-r^5}{1-r} \right)\] Your series starts at 1000, so a = 1000

OpenStudy (anonymous):

\[\huge S_5 = 1000\left( \frac{1-r^5}{1-r} \right)\] And your common ratio is 1/2 = 0.5

OpenStudy (anonymous):

\[\huge S_5 = 1000\left( \frac{1-\left(\frac{1}{2}\right)^5}{1-\frac{1}{2}} \right)\] Now do the number-crunching :P

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

I still didnt get my answer from using my calc.

OpenStudy (anonymous):

Try harder :D

OpenStudy (anonymous):

I got my answer

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

1937.5

OpenStudy (anonymous):

Awesome, you got it! Now, you could have gotten the same answer if you just keyed in 1000+500+250+125+62.5 :D

OpenStudy (anonymous):

lol thank you

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I need more help in the future so stay online ill tag you in the problem i need help with, but thanks a lot.